# Neuron Empirical Gradient: Discovering and Quantifying Neurons’ Global Linear Controllability

### v2

@@ Abstract
Although feed-forward neurons in pre-trained language models (PLMs) can store knowledge and their importance in influencing model outputs has been studied, existing work focuses on finding a limited set of neurons and analyzing their relative importance. However, the global quantitative role of activation values in shaping outputs remains unclear, hindering further advancements in applications like knowledge editing. Our study first investigates the numerical relationship between neuron activations and model output and discovers the global linear relationship between them through neuron interventions on a knowledge probing dataset. We refer to the gradient of this linear relationship as neuron empirical gradient (NEG), and introduce NeurGrad, an accurate and efficient method for computing NEG. NeurGrad enables quantitative analysis of all neurons in PLMs, advancing our understanding of neurons’ controllability. Furthermore, we explore NEG’s ability to represent language skills across diverse prompts via skill neuron probing. Experiments on MCEval8k, a multi-choice knowledge benchmark spanning various genres, validate NEG’s representational ability. The data and code are released. [^1] https://anonymous.4open.science/r/NeurGrad
@@ 1 Introduction
Transformer (Vaswani et al., 2017)-based pretrained language models (PLMs) exhibit a remarkable ability to possess human knowledge, drawing significant attention to understand their internal mechanisms. While prior studies have highlighted the crucial role of feed-forward (FF) layers’ neurons in representing diverse knowledge, including factual knowledge (Dai et al., 2022; Yu and Ananiadou, 2024) and general language skills (Wang et al., 2022; Tan et al., 2024), they face several challenges. First, the existing methods primarily focus on ranking neurons to identify important ones (Dai et al., 2022; Meng et al., 2022; Yu and Ananiadou, 2024), yet they lack a direct and global quantitative measurement between neuron activations and model output, limiting their applicability to neuron-level adjustment applications, like knowledge editing Zhang et al. (2024). Second, the existing neuron importance-ranking methods are computationally expensive, requiring either multiple activation modification (Dai et al., 2022; Meng et al., 2022; Goldowsky-Dill et al., 2023) or extensive tensor calculation (Yu and Ananiadou, 2024), making inefficient to analyze all neurons on large models across diverse prompts, thereby hindering a global understanding of their roles.
Figure 1: Overview of our contributions: i) observation of the neuron linearity, ii) an efficient method, NeurGrad, to quantify the linearity, iii) skill neuron probing on MCEval8K to confirm NEG captures language skills.
@@
Initially, our study conducts a quantitative analysis to understand how neuron activations determine model outputs (RQ1). We investigate this by gradually modifying activations of randomly sampled neurons and observing the resulting changes in the probabilities of target tokens for correct knowledge (hereafter, output shift), using MyriadLAMA (Zhao et al., 2024), a factual knowledge probing dataset on nine PLMs, including large language models (LLMs) like Llama2-70B (§ 3). Notably, we discover that, within a certain range of activations, shifts in neuron activations (hereafter, activation shifts) have a linear relationship with the output shift. We call and quantify the gradient of this linear relationship as neuron empirical gradient (NEG), enabling quantitative neuron analysis.
@@
Next, we ask: can we precisely adjust PLMs’ output probabilities by shifting neuron activations? (RQ2). As estimating NEG requires extensive inference over PLMs, we first introduce NeurGrad (§ 4), an accurate yet efficient method to calculate NEG for the single neuron, grounded in the empirical observation that computational gradients( 4.1) strongly correlate with NEG magnitudes but weakly with their directions. We validate its performance using MyriadLAMA by measuring their relative relationship to ground-truth NEGs. NeurGrad even shows superiority over baselines, including existing neuron-ranking methods (Dai et al., 2022; Yu and Ananiadou, 2024). Leveraging NeurGrad, we further investigate neuron controllability over multiple neurons through multi-neuron intervention (§ 5). We observe that NEG can be accumulated with multiple neurons, while the correlation diminishes as more neurons are involved or larger activation shifts are applied.
Finally, we investigate whether NEG can represent general language skills associated with diverse prompts rather than specific factual prompts (RQ3). To answer this, we conduct skill neuron probing that identifies neurons associated with language skills using NEGs as inputs. While prior studies conducted skill neuron probing, they focus on using neuron activation to build probers, leaving the representational ability of NEG unexamined (Wang et al., 2022; Song et al., 2024). Furthermore, as they only used multi-choice datasets with limited language skills, we introduce MCEval8K, a multi-choice knowledge evaluation benchmark spanning six genres and 22 tasks for comprehensive LLMs evaluation. Our experiment discovers NEG’s ability to represent diverse language skills, providing a basis for future language skill-oriented neuron-based model adjustments.
@@
  • We confirm that activation shifts are linearly correlated to output shifts within a specific activation shift range, referred to and quantified as neuron empirical gradients. (§ 3)
  • We present NeurGrad, an efficient method for estimating neuron empirical gradients(§ 4), and conduct analysis to deepen the understanding of neuron controllability (§ 5).
  • We find NEG’s ability to express language skills by skill neuron probing (§ 6,§ 7).
  • We built MCEval8K, a multi-choice benchmark spanning six knowledge genres and 22 language-understanding tasks (§ 6.2, § F).
2 Related Work
Neuron-level knowledge attribution methods. Existing studies built the connections between knowledge and neurons by measuring the importance scores of neurons to the model prediction on the target knowledge; some of them rely on causal observations of how knowledge changes with neuron modifications (Meng et al., 2022; Geva et al., 2021; Wang et al., 2024; Chen et al., 2023), while others use extensive tensor calculations to estimate neuron contributions (Geva et al., 2022; Lee et al., 2024; Yu and Ananiadou, 2024). However, these methods are computationally expensive, limiting their scalability for large-scale probing across diverse prompts in LLMs. Moreover, they often capture relative relationships (e.g., neuron rankings) rather than the direct, quantitative relationships between specific neurons and model outputs, restricting neurons’ utility in scenarios requiring precise, such as knowledge editing Zhang et al. (2024) and bias mitigation Gallegos et al. (2024) on LLMs. Although gradient-based approaches (Lundstrom et al., 2022; Dai et al., 2022) seek to integrate gradients through neuron intervention, they also suffer from high computational costs.
Skill neuron probing. Neurons in FF layers show the ability to convey specific skills so that using the neuron activations solely can tackle the language tasks, which these neurons are referred to as skill neurons (Wang et al., 2022; Song et al., 2024). Existing studies found neurons can express semantic skills like sentiment classification (Wang et al., 2022; Song et al., 2024) or complex skills, such as style transfer (Lai et al., 2024) and translation (Tan et al., 2024). However, previous research viewed neuron activations as knowledge indicators, and the representational ability of neuron gradients to language skills is not examined, hindering the application of neuron-level model adjustment.
@@ 3 Neuron Linearity to Model Output
This section empirically answers how neurons in PLMs’ FF layers influence model output by observing the resulting change in output tokens’ probabilities for fine-grained neuron-level interventions.
3.1 Neuron-level Intervention Experiment
@@
Models. To make the analysis result general, we experiment with masked and causal LMs of varied sizes and learning strategies. For masked LMs, we use three BERT (Vaswani et al., 2017; Devlin et al., 2019) models: BERTbase, BERTlarge, and BERTwwm. We construct masked prompts and let the model predict the masked token. For causal LMs, we examine diverse open-source LLMs, including instruction-tuned and pre-trained LLMs. The instruction-tuned LLMs are selected from two families: Llama2 (Touvron et al., 2023) with sizes of 7B, 13B, and 70B, and Phi3-mini (Abdin et al., 2024), a model with 3.8B parameters. The pre-trained LLMs are Llama2 models with 7B and 13B parameters. Following the zero-shot prompt setting in Zhao et al. (2024), we instruct them to generate single-token answers. See § E for model details.
@@
Dataset. We utilize a multi-prompt knowledge probing dataset, MyriadLAMA (Zhao et al., 2024), for neuron intervention. MyriadLAMA offers diverse prompts per fact, reducing the influence of specific linguistic expressions on probing results. We focus on single-token probing, where the target answer is represented by a single token. For each PLM, we randomly sample 1000 prompts from MyriadLAMA, where the model correctly predicts the target token. Due to differences in tokenizers, the probing prompts may vary across PLMs, and the results are not comparable across the PLMs.
Neuron-wise intervention. We conduct a neuron-wise intervention to analyze how activation shift affects model outputs. Specifically, we alter the neuron activations within a range of [-10, 10] with a step size of 0.2 to observe the resulting changes in target token output probabilities. To establish a global observation over all neurons while minimizing computational cost, since assessing a single neuron’s effect on one token for one prompt requires 100 inferences, we conduct neuron interventions on randomly sampled neurons.
@@
Figure 2: Average absolute Corr between activation shifts and output shifts.
@@
Result and Analysis To understand how output shifts respond to neuron activation shifts, we calculate the Pearson correlation (hereafter, Corr) between the activation shift ranges and the output shifts of the correct tokens, considering only the absolute Corr values. The absolute Corr is averaged over 10 prompts randomly sampled from 1000 prompts to reduce the computational cost, each with 1000 randomly sampled neurons given a specific shift range (x-axis). Figure 2 shows that all activation shifts and output shifts have a nearly linear, strong Corr, even with a broad range222$\pm$10 is an appropriate range considering the distribution of activations; see § B.1 for details. of $\pm$10. Figure 2 also indicates that activation shifts tend to show stronger Corr with output shifts at smaller shift ranges, consistent across all models. It indicates that the shifting neuron activation with a specific value can have a predictable result in output shifts.
Noted that as all PLMs show similar behaviors, we consider the BERT PLMs and three instruction-tuned Llama2 LLMs for the following analysis.
@@ 3.2 Neuron Linearity
Based on the above discussion, we ask: do neurons generally exhibit linearity to model output? We first define neurons as possessing linearity if their correlation (Corr) is at least 0.95333Since linearity lacks a strict definition, we use Corr¿=0.95 to indicate a strong linear relationship. within a shift range of $\pm$2 from the observation in Figure 2.
We present a quantitative analysis of the prevalence of neuron linearity across different prompts and Transformer layers. Specifically, we report the ratio of neurons exhibiting linearity from 1000 prompts paired with 100 neurons. [^4] We only chose 200 prompts and 100 neurons for Llama2-70B due to the large model size. The ratios of ‘linear’ neurons in BERTbase/large/wwm models are 0.9565/0.8756/0.9564, respectively, and the ratios for Llama-7B/13B/70B are 0.9387/0.9677/0.9208, indicating that majority of neurons exhibit linearity. We analyze the generality of linear neurons across layers and prompts, revealing their widespread presence (§ A). Furthermore, we term the linearity direction as polarity, where neurons are positive/negative if increasing/decreasing their activations enhances the target output probabilities.
Neuron empirical gradient. We quantify the neurons’ linearity and polarity by the gradient of the linear relationship between activation shift and output shifts, which we term neuron empirical gradient (NEG). We calculate NEG as follows: we fit a zero-intercept linear regression between activation shifts and output shifts acquired through neuron intervention, and the regression coefficient is identified as the NEG for a specific neuron, prompt, and token.
Corr MAE
CG IG LPI NeurGrad CG IG LPI NeurGrad
BERTbase -.8909 .7360 - [^5] We follow code released in Yu and Ananiadou (2024) and only Llama2 LLMs are supported. .9998 6.1e-03 3.0e-03 - 2.6e-05
BERTlarge -.9307 .7167 - .9958 4.6e-03 2.1e-03 - 1.9e-04
BERTwwm -.8914 .8584 - .9989 4.5e-03 2.2e-03 - 2.3e-05
Llama2-7B .3020 .5383 .6469 .8136 1.7e-06 1.2e-06 1.9e-03 1.3e-06
Llama2-13B -.1973 .7261 .1141 .9965 2.1e-06 1.5e-06 4.3e-04 6.0e-08
Llama2-70B .0283 n/a [^6] Due to the high memory cost of IG and LPI, we preclude Llama2-70B experiments on these methods. n/a 1.000 5.9e-07 n/a n/a 2.1e-09
Avg. Runtime (Llama2-7B) 0.149s 19.349s 6.086s 0.161s Same as left
Table 1: Evaluation of NeurGrad and baselines in calculating NEGs, including two metrics: Corr and MAE.
@@ 4 NeurGrad for NEG Estimation
Efficiently and accurately computing NEG is crucial for quantitative analysis of neuron-level interpretability in PLMs. However, quantifying NEG through neuron-wise intervention is impractical due to the high computational cost. While prior studies proposed knowledge attribution methods that measure neurons’ importance in influencing model output, these methods require either extensive calculation or can only estimate the neurons’ relative importance but cannot directly estimate the NEG (Dai et al., 2022; Geva et al., 2022; Meng et al., 2022; Yu and Ananiadou, 2024).
@@ 4.1 NeurGrad
where $\bar{G_{E}}$, $A$, and $\operatorname{sign}(A)$ represents the estimated NEG, activation, and sign of $A$ (1 for $A>0$ and -1 for $A<0$), respectively.
@@ 4.2 NEG Estimation Evaluation
We evaluate NeurGrad’s effectiveness in estimating NEG. We use the same setting as CG evaluation above to collect the ground-truth NEGs. Our experiment involves three baselines, including two gradient estimation methods, CG and integrated gradients (IG) (Dai et al., 2022), and one logit-based knowledge attribution method (Yu and Ananiadou, 2024). IG intervenes with neurons in small step sizes multiple times to simulate the NEG. The Log-Probability-Increase (LPI) (Yu and Ananiadou, 2024) identifies important neurons by estimating increased probabilities with specific neurons. [^8] We did not consider other causal-tracing-based methods (Meng et al., 2022) due to their high computational cost, which conflicts with our efficiency goal.
We evaluate how well these methods calculate NEG with two metrics: Corr and mean absolute error (MAE). Table 1 (left) reports the Corr between the estimated and ground-truth NEGs that can assess their ability to capture the neurons’ relative relationship, such as ordering. NeurGrad consistently achieves high Corr over the six PLMs, outperforming all other methods. Moreover, Table 1 (right) reports the Corr the MAE that quantifies the accuracy of NEGs’ calculation. It is observed that NeurGrad can largely reduce the estimation error to the true NEG, revealing its ability to precisely calculate NEGs. Finally, the average running time (Table 1 (bottom)) demonstrates NeurGrad’s efficiency compared to other methods. The time estimation uses Llama2-7B with an NVIDIA RTX A6000 GPU.

### v3

@@ Abstract
While feed-forward neurons in pre-trained language models (PLMs) can encode knowledge, past research targeted a small subset of neurons that heavily influence outputs. This leaves the broader role of neuron activations unclear, limiting progress in areas like knowledge editing. We uncover a global linear relationship between neuron activations and outputs using neuron interventions on a knowledge probing dataset. The gradient of this linear relationship, which we call the neuron empirical gradient (NEG), captures how changes in activations affect predictions. To compute NEG efficiently, we propose NeurGrad, enabling large-scale analysis of neuron behavior in PLMs. We also show that NEG effectively captures language skills across diverse prompts through skill neuron probing. Experiments on MCEval8k, a multi-genre multiple-choice knowledge benchmark, support NEG’s ability to represent model knowledge. Further analysis highlights the key properties of NEG-based skill representation: efficiency, robustness, flexibility, and interdependency. The code and data are released.
[…]
xzhao-tkl/NEG     [Uncaptioned image]iszhaoxin/MCEval8K
@@ 1 Introduction
Pre-trained language models (PLMs) based on Transformer architecture Vaswani et al. (2017) effectively encode human knowledge, prompting efforts to understand their inner workings. While previous studies have shown that the feed-forward (FF) neurons play key roles in encoding factual knowledge (Dai et al., 2022; Yu and Ananiadou, 2024) and general language skills (Wang et al., 2022; Tan et al., 2024), they face two main challenges. First, current methods mostly rank neurons by importance without measuring the link between neuron activations and model outputs (Dai et al., 2022; Meng et al., 2022; Yu and Ananiadou, 2024), limiting use cases like knowledge editing Zhang et al. (2024). Second, these methods are costly, involving repeated activation modifications (Dai et al., 2022; Meng et al., 2022; Goldowsky-Dill et al., 2023) or extensive tensor operations (Yu and Ananiadou, 2024), making them inefficient for analyzing all neurons on large models.
Figure 1: Our contributions: i) revealing linearity between activation and output shifts, ii) proposing NeurGrad, an efficient method to quantify it, iii) confirming that NEGs capture language skills on MCEval8K.
@@
This study begins by quantitatively analyzing how neuron activations influence model outputs (RQ1) (Figrue 1). Using MyriadLAMA (Zhao et al., 2024), a factual knowledge probing dataset, on seven PLMs, including large language models (LLMs) like Llama2-70B (§ 2), we gradually modify random neuron activations and observe the changes in the token probabilities for correct knowledge (hereafter, output shifts). Notably, we find that within a certain range, shifts in neuron activations (hereafter, activation shifts) have a linear impact on output shifts. We define and quantify the gradient of this linear relationship as neuron empirical gradient (NEG), enabling quantitative neuron analysis.
@@
Next, we explore whether shifting neuron activations can precisely control PLMs’ output probabilities (RQ2). Since computing NEGs requires costly inference, we propose NeurGrad (§ 3), an efficient method for estimating a single neuron’s NEG. It builds on the empirical finding that computational gradients (§ 3.1) strongly correlate with NEG magnitudes, though less so with directions. Using MyriadLAMA, we validate NeurGrad’s performance against ground-truth NEGs, showing it outperforms existing neuron-ranking methods (Dai et al., 2022; Yu and Ananiadou, 2024). With NeurGrad, we further examine multi-neuron control, finding that NEGs can accumulate across neurons (§ 4), while the effect weakens as more neurons are involved or larger activation shifts are involved.
Finally, we examine whether NEGs can capture general language skills beyond factual knowledge (RQ3). To explore this, we perform skill neuron probing (Wang et al., 2022; Song et al., 2024) using NEGs. As the previous studies on skill neuron probing focus on limited language skills, we introduce MCEval8K, a benchmark covering six genres and 22 tasks for broad LLM evaluation. Our results show that NEGs can represent a wide range of language skills. Furthermore, our in-depth analysis highlights the properties of NEG-based skill neurons, including the efficiency in representing language skills, robustness when facing diverse contexts, substitutability of skills being represented by different neurons, and interdependency between neurons in representing skills.
@@
  • We confirm that activation and output shifts are linearly correlated within a certain range through neuron intervention, defined as neuron empirical gradient (NEG) (§ 2).
  • We introduce NeurGrad, an efficient method to NEG (§ 3), and conduct in-depth analysis about neuron controllability (§ 4).
  • We show that NEGs can represent language skills through skill neuron probing (§ 5); skill neurons exhibit efficiency, robustness, inclusivity, and interdependency (§ 6).
  • We develop MCEval8K, a multiple-choice benchmark covering six genres and 22 language understanding tasks (§ 5.2, § F).
@@ 2 Neuron Linearity to Model Output
This section empirically answers how neurons in PLMs’ FF layers influence model outputs. We observe the resulting change in output tokens’ probabilities for fine-grained neuron-level interventions.
2.1 Neuron-level Intervention Experiments
Models.
@@ 2.1 Neuron-level Intervention Experiments
To ensure the generality of our findings, we evaluate both masked and causal LMs with varying sizes and learning strategies. For masked LMs, we use two BERT Devlin et al. (2019) models, BERTbase and BERTlarge, and have them predict masked tokens. For causal LMs, we examine five LLMs with different model sizes and language families, including Llama2 (Touvron et al., 2023) (7B, 70B), Llama3.1 (8B), Llama3.2 (3B) (Grattafiori et al., 2024), and Qwen2.5 (7B) (Qwen et al., 2025). All of these models are instruction-tuned. Following Zhao et al. (2024), we use a zero-shot prompt to generate single-token answers. See § E for details.
Dataset.
@@
We use MyriadLAMA (Zhao et al., 2024), a multi-prompt knowledge probing dataset for neuron intervention. Its diverse prompts help reduce bias from specific phrasing. We focus on single-token probing, where the target answer is a single token. For each PLM, we randomly sample 1000 prompts from MyriadLAMA, where the model predicts the token representing the correct answer. Due to tokenizer and knowledge differences, the sampled prompts vary across PLMs.
Neuron intervention.
We shift activations in the range of [-10, 10] [^1] The range mostly covers the distribution of activations; see § B.1 for details. with the step size of 0.2 to track changes in target token probabilities. For example, consider the following prompt:
Predict the [MASK] in each sentence in one word.
Q: [MASK] is the capital of Japan.
A:
We modify the activations of specific neurons and observe how the probability of the target word Tokyo changes at the final position. Since conducting one intervention experiment for a neuron-token-prompt combination requires 100 inferences due to the shift values, we randomly sample [^2] The actual number of randomly sampled neurons will be given in each experiment. neurons to make computation tractable while ensuring broad coverage.
@@
Figure 2: Average absolute correlation between activation and output shifts.
@@
Results. To understand how output shifts respond to neuron activation shifts, we compute the Pearson correlation coefficient ($r$) between activation shifts and output shifts of correct tokens, using absolute $r$ values. To reduce computation, we average $r$ over 10 randomly sampled prompts from the overall 1000 prompts, each with 1000 randomly sampled neurons at specific shift ranges (x-axis). Figure 2 shows a strong, nearly linear correlation across a wide $\pm$10 range, with stronger correlation at smaller shift ranges consistent across all models. This suggests predictable output changes from specific activation shifts. Since all PLMs behave similarly, we focus on BERT models and Llama2 LLMs in subsequent analysis.
@@ 2.2 Neuron Linearity
Based on the findings above, we ask whether neurons generally show linearity with model output. We define neurons as linear if their correlation ($r$) is at least 0.95333We assume $r\geq 0.95$ to be a strong linear relationship. within a $\pm$2 shift range, as observed in Figure 2.
Then, we quantitatively analyze neuron linearity across prompts and Transformer layers by measuring the ratio of linear neurons from 1000 prompts and 100 randomly sampled neurons.  [^4] We only chose 200 prompts and 100 neurons for Llama2-70B due to the large model size. The ratios of linear neurons are reported in Table 1, showing that most neurons in LLMs exhibit linearity. This linearity is widespread across layers and prompts (§ A). We also define polarity as follows: neurons are positive if increasing activations boost target token probabilities, and negative otherwise. The analysis of neuron polarity is deferred to § 4.1.
Linear Positive Negative
BERTbase .9201 .4981 .5004
BERTlarge .9209 .4938 .4981
Llama3.2-3B .9483 .5000 .5000
Llama2-7B .9659 .5157 .4843
Qwen2.5-7B .6549 .5116 .4884
Llama3.1-8B .9540 .5048 .4952
Llama2-70B .9208 .4962 .5039
Table 1: Ratios of linear neurons and neuron polarity over 1000 prompts with 100 neurons.
Neuron empirical gradient.
We quantify neuron linearity and polarity using the gradient of the linear relationship between activation and output shifts, termed neuron empirical gradient (NEG). To compute NEGs, we fit a zero-intercept linear regression between activation shifts and output shifts acquired through neuron intervention; the regression coefficient serves as the NEG for each neuron, prompt, and token.
@@ 3 NeurGrad for NEG Estimation
Efficient and accurate computation of NEG is essential for quantitative neuron-level interpretability in PLMs. However, the neuron-wise intervention is costly. While prior knowledge attribution methods estimate neurons’ influence on model outputs, they either require intensive computation or only provide relative importance, without directly measuring NEG (Dai et al., 2022; Geva et al., 2022; Meng et al., 2022; Yu and Ananiadou, 2024).
@@ 3.1 NeurGrad
where $\bar{G_{E}}$, $A$, and $\operatorname{sign}(A)$ denote the estimated NEG, activation, and sign of $A$ (1 if $A>0$ and -1 if $A<0$), respectively.
Correlation ($r$) MAE
CG IG LPI NeurGrad CG IG NeurGrad
BERTbase -.8909 .7360 - [^6] We follow code released in Yu and Ananiadou (2024) and only Llama2 LLMs are supported. .9998 6.1e-03 3.0e-03 2.6e-05
BERTlarge -.9307 .7167 - .9958 4.6e-03 2.1e-03 1.9e-04
Llama3.2-3B -.1281 .5854 - .8094 3.5e-06 1.8e-06 1.6e-06
Llama2-7B .3023 .5377 .6469 .8135 1.7e-06 1.2e-06 1.3e-06
Qwen2-7B .1043 .2939 - .5862 1.1e-06 7.6e-07 7.7e-07
Llama3.1-8B .0072 .5098 - .7286 4.9e-06 1.9e-06 3.6e-06
Llama2-70B .0283 n/a [^7] Due to the high memory cost of IG and LPI, we preclude Llama2-70B experiments on these methods. n/a 1.000 2.2e-04 n/a 5.9e-07
Avg. Runtime (Llama2-7B) 0.149s 19.349s 6.086s 0.161s Same as left
Table 2: Evaluation of NeurGrad and baselines in calculating NEGs, including two metrics: $r$ and MAE.
@@ 3.2 NEG Estimation Evaluation
We evaluate NeurGrad’s ability to estimate NEG using the same setup as in the CG evaluation. Our experiment compares NeurGrad with three baselines: two gradient-based methods, CG and integrated gradients (IG) Sundararajan et al. (2017); Dai et al. (2022), and one logit-based method (LPI) Yu and Ananiadou (2024). IG simulates NEG by small, repeated neuron interventions, while LPI (Log-Probability-Increase) estimates neuron importance based on increases in output probabilities. [^8] We exclude causal-tracing methods (Meng et al., 2022), which are too costly and do not meet our efficiency goals.
@@
We assess NEG estimation using correlation ($r$) and mean absolute error (MAE) against ground-truth NEG. As show in Table 2, NeurGrad consistently achieves the highest $r$ across the seven PLMs, capturing relative neuron importance well. NeurGrad also yields low MAE, indicating high accuracy. The average running time of Llama2-7B with an NVIDIA RTX A600 GPU (Table 2 (bottom)) demonstrates NeurGrad’s superior efficiency.